OpenStudy (anonymous):

2x^2=6 solve and find the x intercepts of 2x^2-6

OpenStudy (anonymous):

so we are solving: \[2x^2-6 = 0\] First thing we can do is factor out a 2 from both terms: \[2x^2-6=0 \Rightarrow 2(x^2-3) = 0\]

OpenStudy (anonymous):

Dividing by that 2 just leaves us with: \[x^2-3 = 0 \Rightarrow x^2 = 3 \Rightarrow x = \pm \sqrt{3}\] That would be the x -intercepts.

OpenStudy (anonymous):

sqrt of 3?

OpenStudy (anonymous):

OK I see how you did it

OpenStudy (anonymous):

plus or minus the square root of 3.